Evaluate


d^2r/dt^2 at t=0 when


r=(t3+2t)i−3e−2tj+2sin5tk


1
Expert's answer
2021-12-02T09:38:55-0500

"r=(t^3+2t)i−(3e+2t)j+2sin(5t)k\\
\frac{dr}{dt}=(3t^2+2)i-2j+10cos(5t)k\\
\frac{d^2r}{dt^2}=(6t)i-0j-50sin(5t)k\\
=6ti-50sin(5t)k"